A car sounding its horn at 480 Hz is moving with 20 m/s . the driver hears the beats at what frequency. Speed of sound = is 320 m/s. 

Grade:12

1 Answers

Arun
5 years ago
After reflection from the wall, the sound moves towards observer in the car,
γ=(v+v0)/(v-vs)×γ=(340+20)/(340-20)×480= 540 Hz
 
Hope it helps
 
Regards
Arun